Unless stated otherwise, interest is compounded annually, and payments occur at the end of the period. Face value for bonds is $1000.

  1. ASW just reported earnings per share of $2.50 and an annual dividend of $1.40. Their return on equity is 12%.
    1. If investors require 14%, estimate the share price.
    2. Would you expect the PVGO (Present Value of Growth Opportunities) to be positive or negative? Explain.
Answer #1

1)Retention Ratio(b) : 1- (1.40÷ 2.50 )= 44%

Return on equity (r) = 12%

Growth Rate (g) = b× r

g = 0.44× 0.12 = 0.0528 = 5.28%

Share price = D1÷( Ke-g)

P0 = 1.40(1+0.0528) ÷ (0.14 - 0.052)

P0 = $ 16.75

2) PVGO = P0 - ( EPS÷ ke)

PVGO = 16.75 - (2.50 ÷ 0.14)

PVGO = 16.75 - 17.86

PVGO = ($ 1.11)

PVGO is negative. It is better to distribute the earnings rather than reinvest.